APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Thursday January 29, 2009 (MJD 54861) ---=== OBSERVING TEAM ===--- Howard Brewington Kaike Pan Alaina Shelden (training) Steph Snedden (swing) Connie Rockossi (phone support) ---=== OBSERVING PLAN ===--- 29 Jan 2009 54861.0 0.25 BOSS/SEGUE 10.6 hours ---=== OBSERVING SUMMARY ===--- Cart 8 (plate 3215) was declared done. Imaged stripe 69S and the Segue stripe (first good frame 21). See log for more info Strange noise when loading cart 4. We thought it was the slithead doors, but they were open. ---=== OBSERVING LOG ===--- Afternoon: --------- No problems noted during afternoon checkout. Spectro focus was left unchanged. Removed MJD 54833 - 54835 from sdsshost2/sdssfiles1 The disk space is now: /dev/sdb1 917G 836G 35G 97% /data sdssfiles1p:/data1/data 1.5T 834G 559G 60% /files 23:30Z - Started building ventilation. Night: ----- 00:21Z Bias (plate 3198) 00:22Z Bias (plate 3198) 00:35Z Enclosure opened 01:03Z Cart 6 plate 3198 Dec08a_SEGUE2 previous s/n 102.5 01:15Z gotoField used FK5 stars guider frame #0912 01:31Z preCalibs 01:47Z doNextScience 1800 02:25Z Cart 7 plate 3210 Dec08a_SEGUE2 previous s/n 0 02:32Z gotoField used FK5 stars guider frame #1106 02:37Z preCalibs 02:45Z doNextScience 1800 03:18Z doNextScience 1500 03:46Z postCalibs Scan Strip 69S Run 7941 lambda = -173 alt=32 deg 03:52Z install imager 04:13Z run Skippy 04:23Z Everything good, ramp started at frame 12 lambda= -169 04:31Z Science at lambda= -167 04:35Z mu error is off by 10.8" 05:28Z Science until frame 106 lambda= -155 05:30Z End of ramp frame 125 lambda= -152 Segue cluster Stripe run 7942 05:49Z drift started frame 13 We tried to start goDrift, with no success. We called Connie, and she started goDrift remotely from an iop sessio nwithout a tcc connection. We didn't stop the drift, instead just slewed back to the beginning of the field. 05:52Z focus loop running and going back to track frame 18 05:53Z frame 21 is good seeing is 1.1" 06:01Z cluster seen frame 32, dewar 3 06:32Z watcher complained about vs voltage of ccd 55 06:35Z cluster 2 found at frame 90, dewar 3 06:59Z stopping drift frame 130 07:01Z on-telescope bias T bars unlatched run 7943 07:31Z cart 8 plate 3215 Dec08a_SEGUE2 previous s/n 0 07:32Z gotoField, used FK5 stars, cleared offsets, spiral pattern, some searching 07:52Z gotoField, found fibers frame 1319 07:56Z preCalibs 08:05Z doNextScience 1800 08:37Z doNextScience 1800 09:11Z postCalibs 09:18Z doNextScience 1800 09:51Z doNextScience 1800 10:24Z postCalibs 10:33Z Cart 4 plate 3172 Odd clicking noise when installing the cartridge. We thought it was the slithead dorrs, but they were open. Happened on both sides, on the right (when facing the telescope from the loading area) is earlier and louder than on the left. Moved sp1 -2000 because it was a colder temperature and was drifting. 10:40Z gotoField, used FK5 stars frame # 305 10:45Z preCalibs 10:52Z doNextScience 1800 11:25Z doNextScience 1800 11:59Z postCalibs 12:09Z doNextScience 1800 ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 7939 22:24Z 22:55Z 100 O 174.47 -177.82 62 ignore 7940 04:09Z 04:15Z 69 S -171.69 -170.10 25 science 7941 04:20Z 05:32Z 69 S -169.70 -151.69 134 science 7942 05:46Z 07:00Z 100 O -74.77 -56.15 138 science 7943 07:01Z 07:18Z 100 O -55.92 -51.57 43 bias ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- 7940 8 4 0 0.000 0.000 0.000 0.000 -0.00000 0 0 7941 7 4 0 0.000 0.000 0.000 0.000 -0.00000 0 0 7941 56 4 131 10.500 0.147 -1.900 0.154 -0.02557 289 31 7941 109 4 87 9.800 0.147 -1.600 0.132 -0.02594 286 31 7942 16 4 0 0.000 0.000 0.000 0.000 -0.00000 0 0 7942 67 4 0 0.000 0.000 0.000 0.000 -0.00000 0 0 7942 120 4 0 0.000 0.000 0.000 0.000 -0.00000 0 0 ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ 7941 33 1 31 251 21 -2.465 0.317 3.020 0.287 -0.01016 7941 33 1 31 251 29 -1.896 0.191 3.239 0.232 -0.01089 7941 33 1 31 251 25 -1.341 0.261 3.444 0.185 -0.01158 7941 33 1 31 251 29 -0.922 0.228 3.419 0.182 -0.01150 7941 33 1 31 251 29 -0.225 0.431 3.480 0.247 -0.01170 7941 68 1 31 252 27 -2.172 0.258 2.799 0.161 -0.00941 7941 68 1 31 252 20 -1.890 0.194 3.191 0.224 -0.01073 7941 68 1 31 252 25 -1.639 0.165 3.429 0.204 -0.01153 7941 68 1 31 252 22 -1.338 0.421 3.462 0.214 -0.01164 7941 68 1 31 252 27 -0.908 0.336 3.613 0.213 -0.01215 7941 84 1 31 253 29 -1.537 0.247 3.723 0.238 -0.01252 7941 84 1 31 253 31 -1.090 0.288 3.573 0.410 -0.01201 7941 84 1 31 253 35 -0.611 0.337 3.869 0.345 -0.01301 7941 84 1 31 253 27 -0.440 0.530 3.901 0.373 -0.01312 7941 84 1 31 253 19 -0.124 0.495 4.015 0.257 -0.01350 7941 116 1 31 254 23 -1.391 0.125 3.035 0.180 -0.01020 7941 116 1 31 254 22 -0.927 0.197 3.191 0.160 -0.01073 7941 116 1 31 254 31 -0.401 0.350 3.257 0.248 -0.01095 7941 116 1 31 254 34 -0.118 0.151 3.306 0.244 -0.01112 7941 116 1 31 254 36 0.396 0.216 3.505 0.172 -0.01178 7942 22 1 73 327 75 0.942 0.154 -1.232 0.120 0.00414 7942 22 1 73 327 69 0.969 0.167 -1.116 0.131 0.00375 7942 22 1 73 327 75 1.110 0.117 -1.048 0.131 0.00352 7942 22 1 73 327 74 1.176 0.096 -1.051 0.094 0.00353 7942 22 1 73 327 76 1.168 0.110 -1.041 0.110 0.00350 7942 109 1 56 320 73 0.903 0.265 1.525 0.237 -0.00513 7942 109 1 56 320 68 0.714 0.226 1.735 0.194 -0.00583 7942 109 1 56 320 67 0.741 0.218 1.739 0.209 -0.00585 7942 109 1 56 320 79 0.610 0.198 1.834 0.141 -0.00617 7942 109 1 56 320 62 0.551 0.263 1.923 0.143 -0.00646 ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): Y QA Procedures Done (y/n): Y UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 22:50 58513 3.0 arc 22:52 58514 3.0 arc 00:24 58515 0.0 bias ----- sequence 58516, plate -9999 ------- 00:28 58516 0.0 bias ----- sequence 58516, plate 3198 ------- 01:33 58517 10.0 flat 01:36 58518 2.0 arc 02:20 58519 1800.0 target ----- sequence 58520, plate 3210 ------- 02:40 58520 10.0 flat 02:42 58521 2.0 arc 03:17 58522 1800.0 target 03:45 58523 1501.0 target 03:49 58524 10.0 flat 03:51 58525 2.0 arc ----- sequence 58526, plate 3215 ------- 07:59 58526 10.0 flat 08:01 58527 2.0 arc 08:37 58528 1800.0 target 09:10 58529 1800.0 target 09:13 58530 10.0 flat 09:16 58531 2.0 arc 09:51 58532 1801.0 target 10:23 58533 1801.0 target 10:26 58534 10.0 flat 10:29 58535 2.0 arc ----- sequence 58536, plate 3172 ------- 10:48 58536 10.0 flat 10:50 58537 2.0 arc 11:25 58538 1800.0 target 11:58 58539 1801.0 target 12:01 58540 10.0 flat 12:03 58541 2.0 arc 12:41 58542 1800.0 target ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 01:45Z 6 0 -74.33 -0.0032 49.02 -0.0006 123.02 -0.0008 0.999890 02:47Z 7 3210 264.52 -0.0071 70.65 -0.0032 97.09 0.0010 0.999890 08:07Z 8 3215 127.96 0.0005 65.94 -0.0071 -69.23 0.0169 0.999930 11:00Z 4 3172 -19.06 -0.0015 76.52 -0.0018 163.00 0.0350 0.999900 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 01:45Z 6 0 0.99989 1085 2373 290 -74 49.0 -1.3 12 11 -- 1.5" 02:46Z 7 3210 0.99989 1085 2179 150 265 70.8 -1.4 7 19 -- 1.2" 08:07Z 8 3215 0.99993 690 1911 200 128 65.8 -3.1 20 8 -- 1.0" 11:00Z 4 3172 0.99990 986 2039 100 -19 76.5 -0.3 12 23 -- 2.0" ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:49Z 29 0 7 315 (NW) 466 - clear 23:21Z 29 0 5 343 (NNW) 553 - 23:53Z 29 0 12 27 (NNE) 560 - 00:37Z 27 0 13 21 (NNE) 613 - 01:07Z 26 0 11 21 (NNE) 729 - 01:38Z 28 -4 13 41 (NE) 643 - 02:09Z 30 -13 8 357 (NNW) 837 - 02:42Z 29 -11 7 339 (NNW) 714 - 03:14Z 28 -11 5 81 (E) 720 - 03:47Z 27 -6 9 38 (NE) 713 - 04:19Z 27 -13 6 350 (NNW) 735 - clear 04:51Z 28 -11 14 356 (NNW) 1118 - 05:23Z 27 -9 14 355 (NNW) 709 - 05:53Z 27 -8 16 3 (N) 606 - 06:23Z 27 -6 21 359 (NNW) 663 - 06:53Z 27 -5 16 3 (N) 607 - 07:23Z 27 -6 17 356 (NNW) 590 - 07:56Z 27 -4 22 6 (N) 598 - 08:27Z 26 -2 16 5 (N) 585 - 08:59Z 25 -4 11 34 (NE) 552 - 09:32Z 29 -9 16 22 (NNE) 580 - 10:04Z 30 -10 21 27 (NNE) 513 - 10:36Z 31 -10 15 21 (NNE) 416 - 11:08Z 31 -7 9 33 (NNE) 427 - 11:40Z 32 -5 6 45 (NE) 225 - 12:13Z 32 -6 10 47 (NE) 174 - 12:45Z 35 -8 1 133 (SE) 163 - clear ---=== TELESCOPE STATUS ===--- 23:30Z - Started building ventilation 00:35Z Enclosure open 12:45Z Enclosure closed No interlocks bypassed. Status at: 12:58Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 04 Counterweights at: 250 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 278.2lbs, 23.3 psi Imager 164.8lbs, 20.0 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_39_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_43 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.0270 -0.0280 Axial B 0.0480 0.0470 Axial C -0.0360 -0.0360 Trans D -0.7640 -0.7600 Lateral E 10.5918 10.5791 Lateral F 12.1793 12.1793 GALILS Commanded: 10450. -100. 5750. 8600. 10350. 10000. Actual: 10471. -76. 5755. 8616. 10374. 10023. SETMIR VALUES PriDesOrient: 0.00 -11.72 8.60 496.70 640.50 PriOrient: 1.06 -11.65 8.77 495.77 639.59 SECONDARY: ---------- Focus: 0.00 Air Temp.: -1.3 Alt.: 30.034752 MIGS TONIGHT NOMINAL Axial A 18.5840 18.5840 Axial B 15.6190 15.6220 Axial C 13.8150 13.8150 Trans D 15.1060 15.1080 GALILS Commanded: 1626723. 1528758. 1595592. -2950. -4550. Actual: 1626627. 1528808. 1595583. -2945. -4561. SETMIR VALUES SecDesOrient: 1257.00 -12.00 -28.00 -39.00 94.77 SecOrient: 1257.01 -12.00 -28.04 -39.22 94.70 ---=== PROBLEMS IN DETAIL ===--- Daily Spectro Focus Check: -------------------------- Cartridge 6 sp1> Coll_motor_A -5797 01P-0000005797 Coll_motor_B -5996 02P-0000005996 Coll_motor_C -5913 01P-0000005913 Coll_motor_A_status 132 01S0132 Coll_motor_B_status 132 02S0132 Coll_motor_C_status 133 01S0133 Requested_exp.time 0 Exp_time_left 0 Last_exp.time 3 Last_open_transit.time 715 Last_close_transit.time 535 Humid_Hartmann 16.7 0xFDDC Humid_Cen_Optics 13.9 0xFE01 Temp_Median 4.6 Temp_Hartmann_Top 4.2 (0xFF4F) Temp_Red_Cam_Bot 4.6 (0xFF41) Temp_Red_Cam_Top 4.8 (0xFF35) Temp_Blue_Cam_Bot 5.0 (0xFF2D) Temp_Blue_Cam_Top 4.7 (0xFF3C) sp2> Right_open_sensor Off Right_closed_sensor On Coll_motor_A -3938 01P-0000003938 Coll_motor_B -4466 02P-0000004466 Coll_motor_C -4645 01P-0000004645 Coll_motor_A_status 133 01S0133 Coll_motor_B_status 132 02S0132 Coll_motor_C_status 132 01S0132 Requested_exp.time 0 Exp_time_left 0 Last_exp.time 3 Last_open_transit.time 725 Last_close_transit.time 630 Humid_Hartmann -24.6 0x0 Humid_Cen_Optics 15.4 0xFDEC Temp_Median 2.7 Temp_Hartmann_Top 3.1 (0xFF80) Temp_Red_Cam_Bot 3.4 (0xFF70) Temp_Red_Cam_Top 2.5 (0xFF98) Temp_Blue_Cam_Bot 2.8 (0xFF8C) Temp_Blue_Cam_Top 2.7 (0xFF90) AD 23359 1094 The above are good for an air temp of 1 to 2 C. Temps are predicted to be colder than that tonight. However, per Howard's request, I left the actuators alone. The collimate results: Reading /data/spectro/54861/sdR-r1-00058513.fit Reading /data/spectro/54861/sdR-r1-00058514.fit COLLIMATE: Min offset = -0.33 pix COLLIMATE: Max offset = 0.42 pix COLLIMATE: RMS across CCD = 0.21 pix COLLIMATE: Mean offset = -0.03 pix COLLIMATE: Camera r1 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of 238 steps for spectro-1 COLLIMATE: Issue SOP command: "sp1; mechPiston 238" Log file = Collimate-54861-r1-00058513.log PostScript file = Collimate-54861-r1-00058513.ps Reading /data/spectro/54861/sdR-r2-00058513.fit Reading /data/spectro/54861/sdR-r2-00058514.fit COLLIMATE: Min offset = -0.45 pix COLLIMATE: Max offset = 0.41 pix COLLIMATE: RMS across CCD = 0.23 pix COLLIMATE: Mean offset = 0.13 pix COLLIMATE: Camera r2 appears to be IN-FOCUS (|mean| < 0.15 pix) COLLIMATE: Predict (red) piston movement of -1203 steps for spectro-2 COLLIMATE: Issue SOP command: "sp2; mechPiston -1203" Log file = Collimate-54861-r2-00058513.log PostScript file = Collimate-54861-r2-00058513.ps Daily spectro bias check: ------------------------ Filename 02% 05% 10% 50% 90% 95% 98% ------------------- ------ ------ ------ ------ ------ ------ ------ sdR-b1-00058510.fit -5.0 -3.9 -2.8 2.5 7.1 8.8 10.4 sdR-b2-00058510.fit -8.0 -5.5 -4.3 1.9 7.2 9.3 10.9 sdR-r1-00058510.fit -6.6 -5.2 -3.3 1.8 7.6 8.8 10.8 sdR-r2-00058510.fit -7.6 -5.5 -4.4 1.9 7.2 9.3 11.4